Is the equation v = v0 + ax dimensionally correct?


1
Expert's answer
2021-12-22T14:08:16-0500

"v =v_0+a*x"

"\text{Dimension:}"

"v- [\frac{L}{T}]"

"v_0- [\frac{L}{T}]"

"a- [\frac{L}{T^2}]"

"x -[L]"

"[\frac{L}{T}]\not =[\frac{L}{T}]+[\frac{L^2}{T^2}]"

"1\not =1+[\frac{L}{T}]"


"\text{Answer: equation is not dimensionally correct}"